Hartman Motors has $16 million in assets, which were financed with $4.8 million of debt and $11.2 million in equity. Hartman's beta is currently 0.8, and its tax rate is 30%. Use the Hamada equation to find Hartman's unlevered beta, bU. Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Beta Levered =0.8
Tax Rate =30%
Beta Unlevered =Beta Levered /((1+(1-Tax Rate)*Debt/Equity) =0.8/((1+(1-30%)*4.8/11.2) =0.62